# Two undecidable decision problems on an ordered pair of non-negative integers

### v32

@@ Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. Since $f$ is not computable, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))\wedge$ $(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow$ $y_{i}\cdot y_{j}=y_{k}))$. Similarly, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall~(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists~(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
@@
Key words and phrases: computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
@@
3 The first undecidable decision problem on an ordered pair of non-negative integers
@@ 4 A short program in MuPAD that computes $f$ in the limit
5 The second undecidable decision problem on an ordered pair of non-negative integers

### v33

# An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$

@@ Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. We prove that no algorithm takes as input a non-negative integer $n$ and decides whether or not $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$$\wedge$ $(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}~(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
@@
Key words and phrases: arithmetic of $\mathbb{N}$, computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
@@
3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers that concerns the arithmetic of $\mathbb{N}$
@@ 4 A short program in MuPAD that computes $f$ in the limit
5 An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
@@
Corollary 1.
No algorithm takes as input a non-negative integer $n$ and returns the logical value of the following sentence:
$\exists p,q\in\mathbb{N}~((n=2^{p}\cdot 3^{q})\wedge$
$(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
$(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$